OpenStudy (anonymous):

A 30o angle cuts off an arc in a circle with a diameter of 10 in. the length of the arc to the nearest hundredth is _____ in.

OpenStudy (raden):

use this formula : the central angles/360 = length of arc/circumference of circle

OpenStudy (anonymous):

how do I set it up

OpenStudy (raden):

let the length of arc is x, so 30/360 = x/(2pi*d) 30/360 = x/(2*3.14i*10) solve for x

OpenStudy (raden):

oppss... mistake in formula circumference of circle, that should be 30/360 = x/(pi*d) 30/360 = x/(3.14*10) solve for x

OpenStudy (anonymous):

you know how do that on a phone scientific calculator aha

OpenStudy (raden):

u can use the standard only :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer 2.6166666667

OpenStudy (raden):

correct, but too many numbers there :P

OpenStudy (anonymous):

so just 2.62

OpenStudy (raden):

ah, agreed :)

OpenStudy (anonymous):

thank you

OpenStudy (raden):

welcome